# A diverse, overlooked population of Type Ia supernovae exhibiting mid-infrared signatures of delayed circumstellar interaction

### v1

@@ 3.2 Constraints on CSM configuration
The subsequent emission at later times ( yr), in events for which we have well-sampled late-time MIR light curves, have blackbody radii within the ejecta radius together with a slow increase that traces the ejecta radius evolution (SN 2015cp, SN 2017fra, and SN 2017hyn). When combined with the increased dust mass, this suggests the formation of new dust, likely within the cold, dense shell of the ejecta as it interacts with existing CSM, consistent with the formation of new dust reported for SN 2018evt (Wang et al., 2024). While most of the events presented here are largely consistent with each other, SN 2020mvp is a clear outlier. The dust surrounding this object is cooler and more massive than that for the rest of the sample, suggesting a larger separation between the interior shell interacting with the ejecta and the outer shell producing the MIR emission.
@@ 3.3 Limits from spectroscopic and radio follow-up
We obtained follow-up spectroscopy of SN 2020mvp and SN 2022bbt (described in Appendix B) with the goal of identifying broad or intermediate H or He lines typically seen in CSM-interacting SNe (Smith, 2017; Graham et al., 2017, 2019). The left panel of Figure 4 show the spectra extracted at the expected position of the SN. While we detect narrow emission lines from the host galaxy for SN 2020mvp, we do not detect any broad or intermediate emission lines from either source, as seen in late-time spectra of PTF11kx and SN 2015cp. We derive 5 upper limits on the H luminosity of erg s-1 for SN 2020mvp and erg s-1 for SN 2022bbt. To interpret these limits, we compare them to the evolution of the H luminosity for PTF11kx and SN 2015cp, as well as with a SN IIn, known SNe Ia-CSM and other low-luminosity SNe Ia which show delayed CSM interaction via H emission.
The right panel of Figure 4 shows that while the limits on SN 2022bbt are consistent with the fast H decline of SN 2015cp (Graham et al., 2019), the non-detection rules out long-lived H emission such as that seen in PTF11kx (Silverman et al., 2013b; Graham et al., 2017) and in SNe IIn (Tartaglia et al., 2020). Our non-detection of SN 2020mvp rules out SN IIn-like long-lived H emission. We additionally show the luminous MIR emission for these SNe in Figure 4. In contrast to H emission which fades on yr timescales for most interacting SNe Ia, the MIR emission is much longer-lived, lasting thousands of days. Compared to SNe Ia-CSM and SNe Ia showing delayed CSM interaction in the MIR, the population of events of subluminous SNe Ia with late-time H emission (SN 2016jae, SN 2018cqj, and SN 2017fhw) exhibit H luminosities which are fainter at similar phases. If the X-ray or UV radiation powering the H emission is similarly weak in these objects, the corresponding MIR emission would not be detectable, consistent with our non-detections in NEOWISE data (Figure 4).
In our VLA follow-up observations of SN 2020mvp and SN 2022bbt, we did not detect a radio source for either SN, with a 5 flux density limit of Jy at a frequency of 5.5 GHz for both objects. We observed SN 2020mvp at 1522 days post-explosion and SN 2022bbt at 932 days post-explosion. We adopt the formalism of Chomiuk et al. (2016) for and and infer a density of the environment assuming both ISM ( constant) and wind-like () environments. We assume we are in equipartition with shock microphysical parameters, . For SN 2020mvp, we constrain the mass loss rate for a wind-like environment to be M⊙ yr-1 for a wind velocity of km s-1. We constrain the density for an ISM environment to be cm-3. For SN 2022bbt, we constrain the mass loss rate for a wind-like environment to be M⊙ yr-1 for a wind velocity of km s-1. We constrain the density for an ISM environment to be cm-3.
@@
where is the shell ejecta velocity, is the wind velocity, is the mass loss rate, and is an efficiency factor. Using 10 000 km s-1 for the shell ejecta, 10 km s-1 for the wind, and a fiducial value of , we find a mass loss limit of M⊙ yr-1 for SN 2020mvp and M⊙ yr-1 for SN 2022bbt. We note that the efficiency factor is a poorly constrained value that is estimated to be at early times, but eventually drops to zero (Salamanca et al., 1998). We choose since our measurements are at very late times. These limits are consistent with the radio measurements presented above.
@@ 4 Discussion and summary
Our sample of six SNe with MIR signatures of delayed CSM interaction doubles the previously known population, and in particular, substantially increases the diversity of SN Ia sub-types known to exhibit late CSM interaction. We can estimate the fraction of SNe Ia showing late-time MIR brightening by using a sub-sample of our events (SN 2020mvp and SN 2021jun) which were included in the flux-limited ZTF Bright Transient Survey (BTS; Fremling et al. 2020; Perley et al. 2020), which is spectroscopically complete to 18.5 mag. We restrict the BTS sample to SNe classified between 2018 June to 2021 June that pass BTS sample cuts (Perley et al., 2020), to allow for at least 1.5 years of NEOWISE observations after the discovery of the SN. This results in a sample of 1466 SNe Ia, out of which we find two exhibiting late-time MIR brightening. We therefore set a lower limit of 0.1%999This is a lower limit due to our restrictive filtering criteria, which may have excluded interacting SNe Ia that are coincident with complex galaxy backgrounds, active galactic nuclei hosts, or where the brightening was missed in the NEOWISE observing gaps. on the fraction of SNe Ia which exhibit delayed CSM interaction within 1.5 yr of explosion. This is comparable with the rate of SNe Ia-CSM found by Silverman et al. (2013a), Dilday et al. (2012), and Sharma et al. (2023) via searches around peak-light, indicating a similar fraction of SNe Ia exhibiting delayed interaction with CSM shells at cm. This is also consistent with the upper limits on rates of late-onset CSM interaction estimated from GALEX data by Dubay et al. (2022).
The presence of pre-existing dust disfavors the classical double-degenerate progenitor model for SNe Ia, which are usually assumed to occur in clean environments (Maoz et al., 2014; Taubenberger, 2017). Instead, the existence of (potentially multiple) detached CSM shells indicate recent intense mass loss episodes that ended shortly prior to the explosion, which could arise in either a single degenerate scenario (via recurrent novae; Hachisu et al. 1999; Sokoloski et al. 2006; Dilday et al. 2012; Moore & Bildsten 2012; Dimitriadis et al. 2014), or a core-degenerate model with variable mass-loss rates during common-envelope (CE) ejection (Livio & Riess, 2003; Soker et al., 2013; Soker & Bear, 2023). The dust masses inferred from photometry suggest total masses of M⊙ assuming a gas-to-dust mass ratio of . These masses are similar to that estimated for PTF11kx (Dilday et al., 2012), which was argued by Soker et al. (2013) to be too massive to be explained by recent nova shells. Similarly, the delayed dynamical instability model by Han & Podsiadlowski (2006), which invokes dynamical mass transfer at the time of explosion, cannot explain the time-delay between the explosion and the onset of CSM interaction.
@@
In the more-likely core-degenerate case that better explains the CSM masses, slower CE ejection velocities of 10 km s-1 (Soker, 2019) indicate that the CE was ejected yr prior to explosion, consistent with the constraint by Soker (2019) from SN 2015cp (Graham et al., 2019). Our derived rate for late CSM interaction, which we find to be similar to that of SNe Ia showing prompt CSM interaction (SNe Ia-CSM), support the claim in Soker (2022) that the time delay distribution between CE ejection and explosion may be more likely to be constant than the rate that is predicted for the delay time since star formation (Friedmann & Maoz, 2018; Heringer et al., 2019). Soker (2019) predicts that of SNe Ia should experience ejecta–CSM interaction within 30 yr. Applying a flat time delay distribution to this prediction suggests that of SNe Ia show CSM interaction within 1.5 yr, similar to our lower limit of . In another single-degenerate model, Justham (2011) suggests that a rapidly-rotating super-Chandrasekhar WD can remain stable until after its donor’s H-rich envelope is exhausted and contracts or is ejected, leading to a delayed explosion with little-to-no observable interaction at early times, in line with our observations.
@@
Over half of our sample is comprised of overluminous 91T-like and super-Chandra SNe Ia101010PTF11kx and SN 2015cp were also initially classified as SNe Ia-91T at early phases, as are most known SNe Ia-CSM at very early phases (Phillips et al., 2024)., despite their intrinsic rarity—with 91T-like events and super-Chandra events representing only % and % of all SNe Ia respectively (Dimitriadis et al., 2024). Since the overluminous and long-lived nature of their optical light curves suggests that these SNe have progenitors that are more massive than those of typical SNe Ia, their connection to delayed CSM interaction presented here points at a potential double degenerate scenario where the merger and subsequent explosion happens very shortly after the end of the CE phase (Scalzo et al., 2010; Raskin et al., 2014; Siebert et al., 2024). The lower fraction of normal SNe Ia exhibiting delayed CSM interaction then may point to a longer delay from the end of the common envelope to the time of explosion, if they also arise from the double degenerate channel.
@@
The connection between a SN Iax and a potential helium star companion (in SN 2012Z; McCully et al. 2022) suggests that the late interaction seen in the Type Iax SN 2021jun could be with circumstellar He, in a scenario similar to the helium-rich CSM surrounding SN 2020eyj (Kool et al., 2023). While Kool et al. (2023) favored a single-degenerate He star channel for SN 2020eyj, Soker & Bear (2023) suggested that it could also be explained by a core-degenerate scenario where the last common envelope was produced by expansion and subsequent ejection of the He-rich envelope. Future MIR observations with JWST that are sensitive to the total dust mass via MIR continuum measurements may be able to distinguish between these scenarios based on the differing predictions for the total CSM mass in these two cases (Moriya et al., 2019). While our search was aided by the (short) overlap between systematic optical SN searches and the NEOWISE survey, we demonstrate that the luminous and long-lived IR signatures of CSM interaction provide a powerful probe of the environments of SNe Ia—suggesting that ground-based IR surveys like WINTER (Lourie et al., 2020; Frostig et al., 2024) and PRIME (Yama et al., 2023), as well as space-based surveys like NEO-Surveyor (Mainzer et al., 2023) and the Roman space telescope (Rose et al., 2021) are poised to systematically reveal this population.
@@
We would like to acknowledge useful discussions with Melissa Graham, Wenbin Lu, Ken Shen, and Kiyoshi Masui. We thank Kristin Chiboucas for her excellent support with Gemini observations. G. M. acknowledges the support of the National Science Foundation and the LIGO Laboratory. K. D. was supported by NASA through the NASA Hubble Fellowship grant #HST-HF2-51477.001 awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., for NASA, under contract NAS5-26555. K. D. acknowledges support from a MIT Kavli Institute Fellowship.
@@ Appendix A Archival objects and new sample details
We show optical and MIR light curves of PTF11kx and SN 2015cp in Figure 5, and details of our sample filtering in Table 1.
@@
SN 2017fra/ASASSN-17jq was discovered by ASAS-SN and was also observed by ATLAS and Gaia (Prieto & Stanek, 2017). Coulter et al. (2017) spectroscopically classified it as a SN Ia-91T at d. It resides in a late-type galaxy that may be interacting with a companion and is at a redshift of , resulting in a projected offset of 1.3 kpc. Jiang et al. (2021) first identified a connection between a MIR excess in its host and SN 2017fra, reporting single-epoch dust fits. This MIR emission was also identified by Thévenot et al. (2021a), who also noticed that SN 2017fra has a sibling in SN 2016cda (Kelsey, 2024).111111SN 2016cda was the original candidate which appeared in our search, but the temporal coincidence of the MIR emission with SN 2017fra makes it the more likely association. In addition to the MIR data presented in Jiang et al. (2021), we show five additional NEOWISE detection epochs in Figures 1 and 2. The MIR emission rises and fades smoothly over yrs, peaking at within yrs after explosion; SN 2017fra is the most luminous MIR SN Ia (Figures 1 and 2).
@@
SN 2017hyn was discovered by ATLAS and also observed by Gaia (Tonry et al., 2017). It was spectroscopically classified as a Ia-91T by Onori et al. (2017) and is hosted in an inclined spiral galaxy at , for a projected offset of 8.3 kpc. Thévenot et al. (2021a) first noted a NEOWISE detection coincident with the optical peak, which is observed shortly after peak light. The subsequent long-lived MIR emission is similar to that of SN 2017fra, exhibiting a sharp peak at 1.5 years after explosion.
SN 2020mvp
SN 2020mvp was discovered by ATLAS (Tonry et al., 2020), and was also observed by ZTF and Gaia. Its host is a large star-forming late-type galaxy at , resulting in a SN projected offset of 7.6 kpc. SN 2020mvp was classified as an normal SN Ia by ZTF/SEDM (Dahiwale & Fremling, 2020) as part of its Bright Transient Survey, consistent with its optical light curve (Figure 2). The first NEOWISE detection is during the SN photospheric phase, followed by a subsequent fading and brightening episode years later, with remarkably red MIR colors ( mag).
@@
SN 2020yex was discovered by Gaia (Hodgkin et al., 2020) and classified as a Ia-91T by Gromadzki et al. (2020). It resides in a small blue galaxy with no distinct structure at and is at a projected offset of 2.7 kpc. The MIR emission was first identified by Thévenot et al. (2021b), who noted that the W1 luminosity was consistent with that of SNe Ia-CSM. Unlike the other MIR light curves in our sample, the NEOWISE W2 light curve is consistent with a long-lived monotonic decline but the W1 light curve exhibits a peak at yr after the explosion.
@@
SN 2022bbt was discovered by ATLAS (Tonry et al., 2022) and also observed by Pan-STARRS, Gaia, and ZTF in its Bright Transient Survey. It is offset (, 7.9 kpc projected distance at ) from a faint blue host. Tucker (2022) classify it as a SN Ia, while Johansson et al. (2022) classify it as a Ia-pec. Using the spectrum from Tucker (2022) and the supernova fitting code NGSF (Goldwasser et al., 2022), we find evidence for low velocity C II lines, which together with the luminous (peaking at AB mag in -band) light curve, is consistent with SN 2022bbt being a super-Chandrasekhar (09dc-like) SN Ia (Howell et al., 2006). Since this is most recent event in our sample, only one NEOWISE detection is available; however, the luminosity is consistent with the rest of the population (Figure 1).

### v4

@@ 3.2 Constraints on CSM configuration
The subsequent emission at later times ( yr), in events for which we have well-sampled late-time MIR light curves, have blackbody radii within the ejecta radius together with a slow increase that traces the ejecta radius evolution (SN 2015cp, SN 2017fra, and SN 2017hyn). When combined with the potentially increased dust mass, this suggests the formation of new dust, possibly within the cold, dense shell of the ejecta as it interacts with existing CSM, consistent with the formation of new dust reported for SN 2018evt (Wang et al., 2024), though the lack of spectral information prevents the drawing of firm conclusions.
@@ 3.3 Limits from spectroscopic and radio follow-up
We obtained follow-up spectroscopy of SN 2020yex and SN 2022bbt (described in Appendix B) with the goal of identifying broad or intermediate H or He lines typically seen in CSM-interacting SNe (Smith, 2017; Graham et al., 2017, 2019). The left panel of Figure 4 show the spectra extracted at the expected position of the SN. While we detect narrow emission lines from the host galaxy for SN 2020yex, we do not detect any broad or intermediate emission lines from either source, as seen in late-time spectra of PTF11kx and SN 2015cp. We derive 5 upper limits on the H luminosity of erg s-1 for SN 2020yex and erg s-1 for SN 2022bbt (see Appendix B). To interpret these limits, we compare them to the evolution of the H luminosity for PTF11kx and SN 2015cp, as well as with a SN IIn, known SNe Ia-CSM and other low-luminosity SNe Ia which show delayed CSM interaction via H emission.
The right panel of Figure 4 shows that while the limits on SN 2022bbt are consistent with the fast H decline of SN 2015cp (Graham et al., 2019), the non-detection rules out long-lived H emission such as that seen in PTF11kx (Silverman et al., 2013b; Graham et al., 2017) and in SNe IIn (Tartaglia et al., 2020). Our non-detection of SN 2020yex rules out SN IIn-like long-lived H emission. We additionally show the luminous MIR emission for these SNe in Figure 4. In contrast to H emission which fades on yr timescales for most interacting SNe Ia, the MIR emission is much longer-lived, lasting thousands of days. Compared to SNe Ia-CSM and SNe Ia showing delayed CSM interaction in the MIR, the population of events of subluminous SNe Ia with late-time H emission (SN 2016jae, SN 2018cqj, and SN 2017fhw) exhibit H luminosities which are fainter at similar phases. If the X-ray or UV radiation powering the H emission is similarly weak in these objects, the corresponding MIR emission would not be detectable, consistent with our non-detections in NEOWISE data (Figure 4).
In our VLA follow-up observation of SN 2022bbt at 932 days post-explosion, we did not detect a radio source, with a 5 flux density limit of Jy at a frequency of 5.5 GHz. Adopting the formalism of Chomiuk et al. (2016) for , we infer the density of the environment assuming both ISM ( constant) and wind-like () environments. We assume we are in equipartition with shock microphysical parameters with . For SN 2022bbt, we constrain the mass loss rate for a wind-like environment to be M⊙ yr-1 for a wind velocity of km s-1. We constrain the density for an ISM environment to be cm-3.
@@
where is the shell ejecta velocity, is the wind velocity, is the mass loss rate, and is an efficiency factor. Using 10 000 km s-1 for the shell ejecta, 10 km s-1 for the wind, and a fiducial value of , we find a mass loss limit of M⊙ yr-1 for SN 2020yex and M⊙ yr-1 for SN 2022bbt. We note that the efficiency factor is a poorly constrained value that is estimated to be at early times, but eventually drops to zero (Salamanca et al., 1998). We choose since our measurements are at very late times. These limits are consistent with the radio measurements presented above.
@@ 4 Discussion and summary
Our sample of five SNe with MIR signatures of delayed CSM interaction nearly doubles the previously known population, and in particular, substantially increases the diversity of SN Ia sub-types known to exhibit late CSM interaction. We can roughly estimate the fraction of SNe Ia showing late-time MIR brightening by considering SN 2021jun, which was included in the flux-limited ZTF Bright Transient Survey (BTS; spectroscopically complete to 18.5 mag; Fremling et al. 2020; Perley et al. 2020). We restrict the BTS sample to SNe classified between 2018 June to 2021 June that pass BTS sample cuts (Perley et al., 2020), to allow for at least 1.5 years of NEOWISE observations after the discovery of the SN. This results in a sample of 1466 SNe Ia, out of which we find one exhibiting late-time MIR brightening. We therefore set a lower limit of 0.05% on the fraction of SNe Ia which exhibit delayed CSM interaction within 1.5 yr of explosion.101010This is a lower limit due to our restrictive filtering criteria, which may have excluded interacting SNe Ia that are coincident with complex galaxy backgrounds, active galactic nuclei hosts, or where the brightening was missed in the NEOWISE observing gaps. Additionally, SN 2021jun is a member of the underluminous Iax sub-class, indicating a potentially larger rate among SNe Iax compared to the broader SNe Ia population. This is comparable with the rate of SNe Ia-CSM found by Silverman et al. (2013a), Dilday et al. (2012), and Sharma et al. (2023) via searches around peak-light, indicating a similar fraction of SNe Ia exhibiting delayed interaction with CSM shells at cm. This is also consistent with the upper limits on rates of late-onset CSM interaction estimated from GALEX data by Dubay et al. (2022).
The presence of pre-existing dust disfavors the classical double-degenerate progenitor model for SNe Ia, which are usually assumed to occur in clean environments (Maoz et al., 2014; Taubenberger, 2017). Instead, the existence of (potentially multiple) detached CSM shells indicate recent intense mass loss episodes that ended shortly prior to the explosion, which could arise in either a single degenerate scenario (via recurrent novae; Hachisu et al. 1999; Sokoloski et al. 2006; Dilday et al. 2012; Moore & Bildsten 2012; Dimitriadis et al. 2014), or a core-degenerate model with variable mass-loss rates during common-envelope (CE) ejection (Livio & Riess, 2003; Soker et al., 2013; Soker & Bear, 2023). The dust mass lower limits inferred from photometry ( M⊙) suggest lower limits on total CSM masses of M⊙ assuming a gas-to-dust mass ratio of . On the high end, these lower limits on the masses may be similar to that estimated for PTF11kx (Dilday et al., 2012), which was argued by Soker et al. (2013) to be too massive to be explained by recent nova shells, though nova shells may be suitable for situations with lower CSM masses. Similarly, the delayed dynamical instability model by Han & Podsiadlowski (2006), which invokes dynamical mass transfer at the time of explosion, cannot explain the time-delay between the explosion and the onset of CSM interaction.
@@
In the core-degenerate case that may better explain larger CSM masses, slower CE ejection velocities of 10 km s-1 (Soker, 2019) indicate that the CE was ejected yr prior to explosion, consistent with the constraint by Soker (2019) from SN 2015cp (Graham et al., 2019). Our derived rate for late CSM interaction, which we find to be similar to that of SNe Ia showing prompt CSM interaction (SNe Ia-CSM), support the claim in Soker (2022) that the time delay distribution between CE ejection and explosion may be more likely to be constant than the rate that is predicted for the delay time since star formation (Friedmann & Maoz, 2018; Heringer et al., 2019). Soker (2019) predicts that of SNe Ia should experience ejecta–CSM interaction within 30 yr. Applying a flat time delay distribution to this prediction suggests that of SNe Ia show CSM interaction within 1.5 yr, similar to our lower limit of . In another single-degenerate model, Justham (2011) suggests that a rapidly-rotating super-Chandrasekhar WD can remain stable until after its donor’s H-rich envelope is exhausted and contracts or is ejected, leading to a delayed explosion with little-to-no observable interaction at early times, in line with our observations.
@@
Over half of our sample is comprised of overluminous 91T-like and super-Chandra SNe Ia111111PTF11kx and SN 2015cp were also initially classified as SNe Ia-91T at early phases, as are most known SNe Ia-CSM at very early phases (Phillips et al., 2024)., despite their intrinsic rarity—with 91T-like events and super-Chandra events representing only % and % of all SNe Ia respectively (Dimitriadis et al., 2024). Since the overluminous and long-lived nature of their optical light curves suggests that these SNe have progenitors that are more massive than those of typical SNe Ia, their connection to delayed CSM interaction presented here points at a potential double degenerate scenario where the merger and subsequent explosion happens very shortly after the end of the CE phase (Scalzo et al., 2010; Raskin et al., 2014; Siebert et al., 2024). The lower fraction of normal SNe Ia exhibiting delayed CSM interaction then may point to a longer delay from the end of the common envelope to the time of explosion, if they also arise from the double degenerate channel.
@@
The connection between a SN Iax and a potential helium star companion (in SN 2012Z; McCully et al. 2022) suggests that the late interaction seen in the Type Iax SN 2021jun could be with circumstellar He, in a scenario similar to the helium-rich CSM surrounding SN 2020eyj (Kool et al., 2023). While Kool et al. (2023) favored a single-degenerate He star channel for SN 2020eyj, Soker & Bear (2023) suggested that it could also be explained by a core-degenerate scenario where the last common envelope was produced by expansion and subsequent ejection of the He-rich envelope. Future MIR observations with JWST that are sensitive to the total dust mass via MIR continuum measurements may be able to distinguish between these scenarios based on the differing predictions for the total CSM mass in these two cases (Moriya et al., 2019). In addition, our existing observations cannot rule out MIR emission purely from massive dust echoes with particular geometries; future JWST observations can also explore this possibility. While our search was aided by the (short) overlap between systematic optical SN searches and the NEOWISE survey, we demonstrate that the luminous and long-lived IR signatures of CSM interaction provide a powerful probe of the environments of SNe Ia—suggesting that ground-based IR surveys like WINTER (Lourie et al., 2020; Frostig et al., 2024) and PRIME (Yama et al., 2023), as well as space-based surveys like NEO-Surveyor (Mainzer et al., 2023) and the Roman space telescope (Rose et al., 2021) are poised to systematically reveal this population.
@@
We would like to acknowledge useful discussions with Melissa Graham, Wenbin Lu, Ken Shen, Joel Johansson and Kiyoshi Masui. We thank Kristin Chiboucas for her excellent support with Gemini observations. G. M. acknowledges the support of the National Science Foundation and the LIGO Laboratory. K. D. was supported by NASA through the NASA Hubble Fellowship grant #HST-HF2-51477.001 awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., for NASA, under contract NAS5-26555. K. D. acknowledges support from a MIT Kavli Institute Fellowship. D. F.’s contribution to this material is based upon work supported by the National Science Foundation under Award No. AST-2401779.
@@ Appendix A Archival objects and new sample details
We show optical and MIR light curves of PTF11kx and SN 2015cp in Figure 5, and details of our sample filtering in Table 2. We also show NEOWISE W2 reference, stack, and subtraction images in Figure 6.
@@
NEOWISE imaging not from temporally coincident nuclear activity | 1 | 5 |
@@
SN 2017fra/ASASSN-17jq was discovered by ASAS-SN and was also observed by ATLAS and Gaia (Prieto & Stanek, 2017). Coulter et al. (2017) spectroscopically classified it as a SN Ia-91T at d. It resides in a late-type galaxy that may be interacting with a companion and is at a redshift of , resulting in a projected offset of 1.3 kpc. Jiang et al. (2021) first identified a connection between a MIR excess in its host and SN 2017fra, reporting single-epoch dust fits. This MIR emission was also identified by Thévenot et al. (2021a), who also noticed that SN 2017fra has a sibling in SN 2016cda (Kelsey, 2024).121212SN 2016cda was the original candidate which appeared in our search, but the temporal coincidence of the MIR emission with SN 2017fra makes it the more likely association. In addition to the MIR data presented in Jiang et al. (2021), we show five additional NEOWISE detection epochs in Figures 1 and 2. The MIR emission rises and fades smoothly over yrs, peaking at within yrs after explosion; SN 2017fra is the most luminous MIR SN Ia (Figures 1 and 2).
@@
SN 2017hyn was discovered by ATLAS and also observed by Gaia (Tonry et al., 2017). It was spectroscopically classified as a Ia-91T by Onori et al. (2017) and is hosted in an edge-on disk galaxy at , for a projected offset of 8.3 kpc. Thévenot et al. (2021a) first noted a NEOWISE detection coincident with the optical peak, which is observed shortly after peak light. The subsequent long-lived MIR emission is similar to that of SN 2017fra, exhibiting a sharp peak at 1.5 years after explosion.
@@
SN 2020yex was discovered by Gaia (Hodgkin et al., 2020) and classified as a Ia-91T by Gromadzki et al. (2020). It resides in a blue spheroidal galaxy with no distinct structure at and is at a projected offset of 2.7 kpc. The MIR emission was first identified by Thévenot et al. (2021b), who noted that the luminosity was consistent with that of SNe Ia-CSM. Unlike the other MIR light curves in our sample, the NEOWISE light curve is consistent with a long-lived monotonic decline but the light curve exhibits a peak at yr after the explosion.
@@
SN 2022bbt was discovered by ATLAS (Tonry et al., 2022) and also observed by Pan-STARRS, Gaia, and ZTF in its Bright Transient Survey. It is offset (, 7.9 kpc projected distance at ) from a face-on spiral host. Tucker (2022) classify it as a SN Ia, while Johansson et al. (2022) classify it as a Ia-pec. Using the spectrum from Tucker (2022) and the supernova fitting code NGSF (Goldwasser et al., 2022), we find evidence for low velocity C II lines, which together with the luminous (peaking at AB mag in -band) light curve, is consistent with SN 2022bbt being a super-Chandrasekhar (09dc-like) SN Ia (Howell et al., 2006). Since this is most recent event in our sample, only one NEOWISE detection is available; however, the luminosity is consistent with the rest of the population (Figure 1).